Thank you in advance. This is PDE, Fourier series.

Define B: c[0,1]→ R by By = y2(0) + y2(1) Prove or disprove: B is linear operator
Answer #1

solution: B: c[01] R such that . By = y²lo) ty? (1) det Yg (26) = x. Then y, ² x ² cm ny, (o) + y0) = 0+1=1.5 By, let Y2 = x²